OpenStudy (anonymous):

Find the value of x if the perimeter of the parallelogram is 102.4 ft. ( the parallelogram is slanted left and 33 is on the top on the outisde and x and is on the right side outside

OpenStudy (anonymous):



OpenStudy (anonymous):

33+33+x+x=102.4 2x=102.4-66 2x=36.4 x=18.2

OpenStudy (anonymous):

Got it, thanks.